OpenStudy (anonymous):

Write the equation of the line that passes through the points (1, 7) and (5, 15) using function notation y = 2x + 5 y = 4x + 8 f(x) = 2x + 5 f(x) = 4x + 8 

Mehek (mehek14):

slope: \(\dfrac{y_2-y_1}{x_2-x_1}\)

OpenStudy (alexandervonhumboldt2):

use the magic formula called the slope formula: \[\frac{ y_1-y_2 }{ x_1-x_2 } ~or~\frac{ y_2-y_1 }{ x_2-x_1 }\]

OpenStudy (anonymous):

slope is 2?

OpenStudy (alexandervonhumboldt2):

\[substitute ~~~your~~~ points~~\frac{ 7-15 }{ 1-5 }=\frac{ -8 }{ -4 }=2\]

OpenStudy (alexandervonhumboldt2):

yes 2

Mehek (mehek14):

so you have y= 2x + b you need the y-intercept choose any ordered pair

OpenStudy (anonymous):

so C?

OpenStudy (alexandervonhumboldt2):

now we have the y=mx+b where m is slope. m=2. y=2x+b. subsiitute one of your points into the equation and solve: yes

OpenStudy (alexandervonhumboldt2):

yes it is C

Mehek (mehek14):

yes C function notation = f(x)